OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 1)^2 – 16? Identify the x–intercepts by graphing the quadratic equation. a) x–intercepts: (3, 0) and (–5, 0) y–intercept: (0, –16) b) x–intercepts: (–3, 0) and (5, 0) y–intercept: (0, -15) c) x–intercepts: (–3, 0) and (5, 0) y–intercept: (0, –16) d) x–intercepts: (3, 0) and (–5, 0) y–intercept: (0, –15)

OpenStudy (anonymous):

y intercept is what you get when your replace x by 0 it would be \[y=(0-1)^2-16=1-16=-15

OpenStudy (anonymous):

x intercept is what you get when you set y = 0 it would be \[0=(x-1)^2-16\] \[(x-1)^2=16\] \[x-1=4\] or \[x-1=-4\] \[x=5\] or \[x=-3\]

OpenStudy (anonymous):

so y intercept is (0-15) x intercepts are (-3,0) and (5,0)